Problem Solving

BBC Company is closing its books and gathered the following items for its bank reconciliation:

a) Bank statement contains an ending balance of P290,500 on February 28, 2021, whereas the

company’s ledger shows an ending balance of P260,900

b) Bank statement contains a P100 service charge for maintaining the checking account

c) Bank statement contains interest income of P20

d) BBC issued checks of P50,000 that have not yet been cleared by the bank

e) BBC deposited P20,000 but this did not appear on the bank statement

f) BBC issued checks of P30,000 that have been cleared by the bank

g) BBC deposited P40,000 and this deposit appeared on the bank statement

h) A check for the amount of P470 issued to the office supplier was misreported in the cash payments

journal as P370.

i) A note receivable of P9,800 was collected by the bank.

j) A check of P520 deposited by the company has been charged back as NSF.

k) A monthly auto debit loan payment for P10,000 A refund of a previous bank charge P500

l) The bank has erroneously charged ABC company’s check worth P4,500 but was able to correct it

within the day as they balance and close their transactions.


Required:

1. Prepare a bank reconciliation using adjusted balance method.

2. Compute the total credit memo to be reflected on the reconciliation.

3. Compute the total debit memo to be reflected on the reconciliation.

4. Compute the total bank error.

5. Compute the total book error.

6. Compute the total outstanding checks.

7. Compute the total deposits in transit.

8. How much is the adjusted balance per book and per bank.

9. Prepare adjusting entries on the books of BBC company


1
Expert's answer
2021-02-25T18:17:11-0500

2.100+10000+520=10 620

3.20+9800+100+500=10420

4.260 700-290 500=-29 500

5.260 700-260 900=-200

6.two checks

7.50 000

8.

debit cash register credit checking account - 20

debit cash register credit checking account - 9800

debit cash register credit checking account - 100

debit cash register credit checking account - 500


debit current account credit cash register - 100

debit current account credit cash register - 10 000

debit current account credit cash register - 520